find the value of "a" if a line containing the point (a,-3a)has a y-intercept of 7 and a slope of -2/3
Answer by benazir.sj@gmail.com(109)

solution:
Given
(x1,y1)= (a,-3a)
slope m = -2/3
yintercept b=7

the slope point form is
y-y1 =m(x-x1)
substitute the given values of x1 and y1
y-(-3a) = -2/3(x-a)
y+3a =-2/3x+2/3a
y =-2/3x +2/3a-3a
y=-2/3x +2/3a-9/3a
y=-2/3x-7/3a
the equation was in the form of slope intercept
y=mx+b
ie, y=-2/3x-7/3a
and also we know that y intercept b =7
ie,-7/3a =7
-7a =7*3
a=-3